# Improving Student-AI Interaction Through Pedagogical Prompting: An Example in Computer Science Education

### v1

@@ 8.2.2. Meta-cognitive skills can further strengthen students’ ability to manage AI-enhanced learning
8.3. AI Systems: Providing more accessible learning experiences with pedagogical and data-driven features.
@@
8.3.1. Implementing Pedagogical Features
Given that learners are not expected to possess deep knowledge of learning sciences or instructional design, the responsibility for pedagogical alignment should not rest solely on them. Instead, AI systems should increasingly shoulder this burden by embedding pedagogical features directly into the prompting phase. One promising direction is to offer structured prompt templates or pre-configured scaffolds encapsulated pedagogical strategies (e.g., worked examples, reflective questions) that learners can select from with minimal effort (Kazemitabaar et al., 2024; Xiao et al., 2024a). This not only reduces the cognitive barrier to engaging in pedagogical prompting but also supports students with varying levels of self-regulation and AI literacy.
8.3.2. Implementing Data-Driven features
@@
9. Limitations and Future Work
9.1. Toward Longitudinal Evaluation of Pedagogical Prompting
@@
9.2. More Accurate Measurement on Learners’ Illusion of Competence and Confidence Calibration
@@ 10. Conclusion
Appendix A Appendix: Example Implementations of Pedagogical Prompt
A.1. Generating Worked Example
@@
A.2. Generating Guiding Questions
@@
A.3. Generating In-Context Explanation
@@
Appendix B Appendix: Validation Criteria in System Prompts
B.1. Problem Type
B.1.1. Understand the Task Description
@@
B.1.2. Plan the General Idea of the Solution
@@ B.1.3. Write the Code
B.1.4. Fix a Bug with Error Message
@@ B.2. Context Provided
B.3. Learning Method
B.3.1. Contextualized Explanation
@@
B.3.2. Example Code on Similar Problem
@@
B.3.3. Step-by-Step Guiding Questions
@@
B.4. Learner Level
B.4.1. Beginner Python Programmer
@@
B.4.2. Advanced Python Programmer

### v2

@@ 8.2.2 Meta-cognitive skills can further strengthen students’ ability to manage AI-enhanced learning
8.3 AI Systems: Providing more accessible learning experiences with pedagogical and data-driven features.
@@
8.3.1 Implementing Pedagogical Features
Given that learners are not expected to possess deep knowledge of learning sciences or instructional design, the responsibility for pedagogical alignment should not rest solely on them. Instead, AI systems should increasingly shoulder this burden by embedding pedagogical features directly into the prompting phase. One promising direction is to offer structured prompt templates or pre-configured scaffolds encapsulated pedagogical strategies (e.g., worked examples, reflective questions) that learners can select from with minimal effort (Kazemitabaar et al., 2024; Xiao et al., 2024b). This not only reduces the cognitive barrier to engaging in pedagogical prompting but also supports students with varying levels of self-regulation and AI literacy.
8.3.2 Implementing Data-Driven features
@@
9 Limitations and Future Work
9.1 Toward Longitudinal Evaluation of Pedagogical Prompting
@@
9.2 More Accurate Measurement on Learners’ Illusion of Competence and Confidence Calibration
@@ 10 Conclusion
11 Appendix: Example Implementations of Pedagogical Prompt
11.1 Generating Worked Example
@@
11.2 Generating Guiding Questions
@@
11.3 Generating In-Context Explanation
@@
12 Appendix: Validation Criteria in System Prompts
12.1 Problem Type
12.1.1 Understand the Task Description
@@
12.1.2 Plan the General Idea of the Solution
@@ 12.1.3 Write the Code
12.1.4 Fix a Bug with Error Message
@@ 12.2 Context Provided
12.3 Learning Method
12.3.1 Contextualized Explanation
@@
12.3.2 Example Code on Similar Problem
@@
12.3.3 Step-by-Step Guiding Questions
@@
12.4 Learner Level
12.4.1 Beginner Python Programmer
@@
12.4.2 Advanced Python Programmer
@@ 12.5 Do Not Provide
13 Declaration of Generative AI and AI-Assisted Technologies in the Writing Process
13.1 The use of generative AI and AI-assisted technologies in scientific writing
Grammarly, Notion AI, and ChatGPT were used in the paper refinement stage to improve the readability and language of the work. All revisions were made on the word or sentence level under human oversight, and we carefully reviewed and edited every part of the content.
13.2 The use of generative AI and AI-assisted tools in figures, images and artwork
During the user study of this project (prior to the paper writing stage), the author(s) used OpenAI DALL·E3 to generate visual elements (cartoon characters) for the user study materials. These images were incorporated into comics and were then edited by the author(s) to fit the narrative of the programming scenarios. Visual demonstrations of these materials are included in Figure 4 and Figure 5. This use aligns with elsevier GenAI use policy, as the AI-generated visuals were part of the user study and were not used to manipulate or obscure data, or included in the paper writing process. The tool, version, and process (generating and hand-editing comics) are described to ensure transparency and reproducibility. After using this tool/service, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the published article.